# Vertex-Minimal Paper Tori

### v4

@@ 3.1 Main Result
3.2 The Vectors
Logically speaking, it does not matter how we produced the vectors above, but here we describe our method. Define
$\phi(x)=2x^{\prime}-1,\qquad x^{\prime}=x-{\rm floor\/}(x).$ (10)
The range of $\phi$ is [1,1][-1,1]. Next, we define
$V_{n}=\frac{W_{n}}{\|W_{n}\|},\hskip 30.0ptW_{n}=(\phi(n\sqrt{2}),\phi(n\sqrt{3}),\phi(n\sqrt{5})).$ (11)
The countable collection $\{V_{n}\}$ is dense in the unit sphere. For each pair $(\Delta_{0},\Delta_{1})$ of $T$ we try to verify a numerical version of Equation 7 or 9 (as the case may be). We found by trial and error that taking $\lambda=5\times 10^{-3}$ in the no-common-vertex case and $\lambda=10^{-3}$ in the $1$-common vertex case works well. For each pair, the search yielded some $V_{n}$. We then define
$L_{n}={\rm floor\/}(10^{6}V_{n}),$ (12)
This operation is applied componentwise. This gives us our lists above.
4 Near Flatness
4.1 The Main Result
Let $T$ be the torus from Equation 1. We say that $T$ is $\epsilon$-flat if $\max_{a}|\theta_{a}-2\pi|<\epsilon.$ The maximum is taken over all $8$ cone angles. In this chapter we prove the following result using rational arithmetic.
Lemma 4.1 (Near Flat)
$T$ is $10^{-30}$-flat.
Let $\Omega$ be the integer torus we get by scaling $T$ up by $10^{32}$. It suffices to prove that $\Omega$ is $10^{-30}$-flat. We mention that we computed, using $100$ digit precision arithmetic and $100$ digit precision versions of the square-root and arccos functions, that $\Omega$ is $(7.26\times 10^{-32})$-flat. Some readers might accept this calculation as “good enough proof” and skip the rest of this chapter.
Let $\theta_{i}$ denote the cone angle of $\Omega$ at the $(i)$th vertex. Here we discuss the idea behind the rational arithmetic proof of the Near Flat Lemma. Figure 1.3 shows a plot of the intrinsic flat structure on $\Omega$ up to an error of about $10^{-16}$. Around each vertex $V_{i}$ we have $6$ triangles $T_{i0},...,T_{i5}$. These triangles do not necessarily fit perfectly in the plane, but the error is so small that you cannot tell from the naked eye. Call the union of these $6$ triangles a flower. The $(i)$th flower corresponds to the $(i)$th star in $\Omega$, though only approximately.
For our proof below, we compute high precision versions of the flowers, then scale the coordinates up by $10^{32}$, then round down all the coordinate to integers. This gives us integer models which approximate the geometry of the $8$ stars in $\Omega$. Now something very nice happens. For our scaled-up models, these big integer triangles fit together perfectly around a vertex. Put another way, for each flower $F$ there are $6$ integer vectors $V_{0},...,V_{5}$ such that each of the $6$ triangles in $F$ is defined by $2$ cyclically consecutive vectors on the list. We then compute the mismatch between the various dot products associated to $\Omega$ and the corresponding dot products associated to the integer flowers, and this gives us the Near Flat Lemma. We list the coordinates for our flowers at the end of the chapter.
Rather than compare the angles, we compare the dot products which go into the calculation of the angles. We finish off the argument by using the Lipschitz properties of the function that converts between the relevant dot product ratio and the angle. We perform exact integer arithmetic calculations using the BigInteger class. Out of a slight laziness, and a desire for a transparent argument over a tedious exercise that anyone can perform, we will avail ourselves of the BigDecimal class in Java, which is perfectly capable of computing the ratios of $64$ digit integers up to $100$ digits of precision.
4.2 The Proof
Let $\tau_{ab}$ be the $(b)$th triangle in the $(a)$th star of $\Omega$ and let $\tau_{ab}^{*}$ be the $(b)$th triangle in the $(a)$th integer flower. Let $V_{ab}$ and $W_{ab}$ be the two vectors that $\tau_{ab}$ defines. By this we mean that $V_{ab}$ points from vertex $a$ to the first other vertex of $\tau_{ab}$ and $W_{ab}$ points from vertex $a$ to the second other vertex of $\tau_{ab}$. (The torus is oriented, so we have no trouble defining “first” and “second” here.) We define $V_{ab}^{*}$ and $W_{ab}^{*}$ with respect to $\tau_{ab}^{*}$ in a similar manner. Let $\theta_{ab}$ and $\theta^{*}_{ab}$ respectively denote the angles of $\tau_{ab}$ and $\tau_{ab}^{*}$ at vertex $(a)$.
When $r>0$ we have
$\theta=\phi(r),\hskip 30.0pt\phi(\cdot)=\arccos(\sqrt{(\cdot)}),\qquad r=\frac{(V\cdot W)(V\cdot W)}{(V\cdot V)(W\cdot W)},$ (13)
We take the positive branch of the square root function here.
We have $48=8\times 6$ pairs of triangles to check. We get the following bounds which work uniformly for all these pairs:
  1. 1.
    $r,r^{*}\in I:=[.00112,.77714]$ for all $a,b$.
  2. 2.
    $|r-r^{*}|<1.07\times 10^{-32}$.
  3. 3.
    $V\cdot W>0$ and $V^{*}\cdot W^{*}>0$.
Lemma 4.2
$\phi$ is $15$-lipschitz on $I$.
Proof: We compute that
$\bigg{|}\frac{d\phi}{dt}\bigg{|}=\frac{1}{2\sqrt{t(1-t)}}.$
Establishing our bound is the same as establishing that
$225-\frac{1}{4t(1-t)}>0,\hskip 30.0pt\forall t\in I$
Establishing this bound is a routine exercise in algebra (or calculus) which we omit. $\spadesuit$
Combining Lemma 4.2 with our calculations, and remembering that our flowers have $6$ triangles in them, we see that $\Omega$ is $\epsilon$-flat when
$\epsilon=6\times 15\times 1.07\times 10^{-32}<10^{-30}.$
This completes the proof of the Near Flat Lemma.
4.3 The Integer Flowers
$\begin{bmatrix}124788644050109876602825361295674&0&0\\ 77712297261878079581593486900079&70129923825301614779354253224291&0\\ -70435981294870011898311359845207&186386084617464728269924163644524&0\\ -148148278556748091479904846745286&116256160792163113490569910420232&0\\ -98095839650912348417428186459704&-84460765042234115458262663828023&0\\ 53415963724939232100166950154753&-89151222668009434223254123690615&0\end{bmatrix}$
$\begin{bmatrix}84465310291672221527450441188266&0&0\\ 69550368811175622953657281341642&103609612884589392577024128236871&0\\ -34241316280724259280029981722340&108947382985381561571026048979621&0\\ -103791685091899882233687263063983&5337770100792168994001920742749&0\\ -88338793860668172702704004953884&-123183133805664642986337651698562&0\\ 44045801850979783812535215764631&-101107834979521394561120290119996&0\end{bmatrix}$
$\begin{bmatrix}104677635440723967179988723412878&0&0\\ 12034998257537687883214969911430&83603513440581754781321562053477&0\\ -94317081172076477932017348008492&54413409707656384418499405421507&0\\ -191630826036432582723537220778797&-128131286236400952463466737043786&0\\ -97313744864356104791519872770305&-182544695944057336881966142465293&0\\ 32097662001199218957367514938600&-185561707945141208999189687748482&0\end{bmatrix}$
$\begin{bmatrix}151584388889297911866446233050267&0&0\\ 56495441368939829023723230650622&94715849040541184436182819641071&0\\ -49223826337965360195502746428317&68639664650392174112843586146043&0\\ -124728889560435721214027036255644&-3861320096243261090662872882149&0\\ -68579908206882346371371696550992&-91317009315546418606880471182019&0\\ 56148981353553374842655339704651&-87455689219303157516217598299871&0\end{bmatrix}$
$\begin{bmatrix}199251097863976649156276288186992&0&0\\ 161120818626134715906111075736407&97485840852430004244786950552576&0\\ -44330167652533400164597159444753&121619286399136814961872050344316&0\\ -102277745880424119662590582022408&-18451785319703080875434679167189&0\\ -44113235471988020751569542690915&-116731435954750256516337414942322&0\\ 38130279237841933250165212450584&-97485840852430004244786950552577&0\end{bmatrix}$
$\begin{bmatrix}206863556376741568622946129706945&0&0\\ 3647481315941312863078263678069&108826590670126188269802629065616&0\\ -75789172382575842209671187858145&80117590481155307070115264260041&0\\ -116999860561305386939755847152992&-65757376990228984547591251604189&0\\ -41210688178729544730084659294848&-145874967471384291617706515864231&0\\ 58216001420386574195913238485630&-115617091154199741307099293537955&0\end{bmatrix}$
$\begin{bmatrix}188317304997339823174158679929020&0&0\\ 17841734520938330693404444524334&103145915434133734843279705318812&0\\ -80328865359490685097501135729666&26108658194654254413402678301333&0\\ -69693027059078945150953187346448&-83662494202455064363150318793564&0\\ 93593941525210768327841471388803&-210666025048595406864753841069280&0\\ 163286968584289713478794658735251&-127003530846140342501603522275717&0\end{bmatrix}$
$\begin{bmatrix}151584388889297911866446233050267&0&0\\ 95435407535744537023790893345615&87455689219303157516217598299870&0\\ -56239231360498057885401519560362&199072046789130364493850160222280&0\\ -105719267706905189219225977078940&-26076184390149010323339233495028&0\\ -56495441368939829023723230650623&-94715849040541184436182819641072&0\\ 95088947520358082842723002399644&-94715849040541184436182819641072&0\end{bmatrix}$
5 Proof of Theorem 1.2
5.1 A Topological Result
Throughout this chapter $\|\cdot\|$ denotes the Euclidean norm and $\|\cdot\|_{\infty}$ denotes the maximum of the absolute values of vector or matrix entries.
@@ 5.1 A Topological Result
Let $G:\mbox{\boldmath{$R$}}^{3}\to\mbox{\boldmath{$R$}}^{3}$ be a smooth map. Let $B_{r}(p)$ denote the ball of radius $r$ about $p$. We say that $G$ is $\lambda$-expansive on $B_{r}(p)$ if, for all unit vectors $V$ and all $q\in B_{r}(p)$, we have the following conditions.
  • The angle between $dG_{p}(V)$ and $dG_{q}(V)$ is less than $\pi/3$.
  • $\|dG|_{q}(V)\|>2\lambda$.
Here we are interpreting $V$ as a vector in all tangent spaces simultaneously.
@@
Proof: Let $O$ be the origin and let $B=B_{1}(O)$. We can scale and translate so that $\lambda=1$ and $p=q=O$. We need to prove that $B\subset G(B)$.
Let $\alpha$ be any line segment connecting $O$ to a point on $\partial B$. Let $V_{\alpha}$ be the unit vector which coincides with $\alpha$. Our hypotheses on $G$ guarantees that the projection of $G(\alpha)$ onto $V_{\alpha}$ always moves with speed greater than $1$. Hence $G(\alpha\cap\partial B)$ lies outside $B$. Given the angle condition, the sphere $G(\partial B)$ is the generator of the homology group $H_{2}(\mbox{\boldmath{$R$}}^{3}-q)$ for any $q\in B$. This easily implies that $B\subset G(B)$. $\spadesuit$
5.2 The Proof Modulo a Bound
Let $F:(z_{0},z_{1},z_{2})=(\theta_{0},\theta_{1},\theta_{2})$ be the map from §1.5. Let $p\in\mbox{\boldmath{$R$}}^{3}$ be the point that corresponds to our torus from Equation 1.1. For the rest of the chapter let $B=B_{10^{-28}}(p)$. Below we will establish the following bound.
@@ 5.2 The Proof Modulo a Bound
Throughout $B$, we have
$\bigg{|}\frac{\partial^{2}F_{k}}{\partial z_{i}\partial z_{j}}\bigg{|}<10^{25},\hskip 30.0pt\forall\ i,j,k\in\{0,1,2\}.$
Now we have all the ingredients needed for the proof of Theorem 1.2. Define
$M=\left[\begin{matrix}-0.91&+0.74&+0.39\\ +0.74&-1.92&+1.14\\ +0.39&+1.14&-0.06\end{matrix}\right]$ (14)
Lemma 5.3
$\|dF_{q}-M\|<1/50$ for all $q\in B$.
Proof: The equation for the Jacobian $dF_{p}$, given in Equation 3, tells us that
$\|M-dF_{p}\|_{\infty}<\frac{1}{100}.$ (15)
Integrating the bound from the Crude Bound Lemma along a rectilinear path of length at most $\sqrt{3}\times 10^{-28}$ that joins $p$ to $q$ we see that
$\|dF_{q}-dF_{p}\|_{\infty}<\sqrt{3}\times 10^{-28}\times 10^{25}<\frac{1}{100}.$
This combines with Equation 15 and the triangle inequality to prove this lemma. $\spadesuit$
@@
$F$ is $(1/10)$-expansive on $B$.
Proof: Let $V$ be any unit vector. The eigenvalues of $M$ all exceed $0.76$ in absolute value. Since $M$ is a symmetric matrix, we have $\|M(V)\|\geq 3/4$.
By the previous lemma we can write $dF_{q}=M+E_{q}$ where $\|E_{q}\|_{\infty}<1/50$. Note that $\|E_{q}(V)\|<3/50$. Since $\|M(V)\|>3/4$ and $\|E_{q}(V)\|<3/50$, we have $\|E_{q}(V)\|/\|M(V)\|<2/25$, which easily implies that the angle between $M(V)$ and $M(V)+E_{q}(V)$ is less than $\pi/3$. We also have
$\|dF_{q}(V)\|\geq\|M(V)\|-\|E_{q}(V)\|\geq\frac{3}{4}-\frac{3}{50}>\frac{1}{5}.$
Hence $F$ is $(1/10)$-expansive in $B$. $\spadesuit$
Recall that our torus $T$ corresponds to the point $p$. Let $q=F(p)\in\mbox{\boldmath{$R$}}^{3}$ and also let $q^{*}=(2\pi,2\pi,2\pi)$. Since $T$ is $10^{-30}$-flat, we have $\|q-q^{*}\|_{\infty}<10^{-30}$. Hence $q^{*}\in B_{10^{-29}}(q)$. Combining this with the Expansion and Surjectivity Lemmas, we have
$q^{*}\in B_{10^{-29}}(q)\subset F(B_{10^{-28}}(p)).$ (16)
@@
Hence there is some $p^{*}\in B_{10^{-28}}(p)$ such that $F(p^{*})=q^{*}$. But then the torus $T^{*}$ corresponding to $p^{*}$ is flat. Since $T$ is $10^{-4}$-robustly embedded, we see that $T^{*}$ is also embedded. This completes the proof of Theorem 1.2 modulo the Crude Bound Lemma.
5.3 Breaking Down the Terms
Before getting to the proof of the Crude Bound Lemma, I want to discuss a certain robustness in our argument because I don’t want our endgame to give the impression that somehow everything boils down to finely tuned estimates. We have arranged things so that we need a bound of $10^{25}$ on some partial derivatives. By working with more precision, we could have reduced the problem to needing a bound of, say, $10^{1000}$. Our $32$-digit example simultaneously gives us something that does not require us to sweat too much over the final estimate and something which lets us have relatively short data files.
Now we get to the details. Let $F(z_{0},z_{1},z_{2})=(\theta_{0},\theta_{1},\theta_{2})$ be the map we have been studying. We write the $(i,j)$th entry in the Jacobian $dF$ as $\partial_{i}\theta_{j}$. We are considering symmetric variations. We do not just move $(z_{0},z_{1},z_{2})$ but also we move $(z_{4},z_{3},z_{6})$ in a symmetric way. Thus, for instance
$\partial_{0}\theta_{j}=\frac{\partial\theta_{j}}{\partial z_{0}}+\frac{\partial\theta_{j}}{\partial z_{4}}.$
When we take second derivatives we get $4$-term expressions. For instance
$\partial_{01}\theta_{j}=\frac{\partial^{2}\theta_{j}}{\partial z_{0}z_{1}}+\frac{\partial^{2}\theta_{j}}{\partial z_{4}z_{1}}+\frac{\partial^{2}\theta_{j}}{\partial z_{0}z_{3}}+\frac{\partial^{2}\theta_{j}}{\partial z_{4}z_{3}}$
@@ 5.3 Breaking Down the Terms
Each $\theta_{j}$ is a $6$-term expression of terms of the form $\theta_{ijk}$, where $\theta_{ijk}$ is the angle of triangle $(i,j,k)$ at vertex $i$. Hence, $\partial_{ij}\theta_{j}$ is a $24$-term sum of expressions of the form
$\partial(a,b,c,d,e)=\frac{\partial^{2}\theta_{cde}}{\partial z_{a}\partial z_{b}}.$ (17)
These expressions vanish when $\{a,b\}\not\subset\{c,d,e\}$.
We call $\partial(a,b,c,d,e)$ friendly if $c\not\in\{a,b\}$. Otherwise we call the term unfriendly. Our terminology indicates our preference.
@@
Proof: First consider terms of the form $\partial(a,a,a,b,c)$. Using the relation $\theta_{a}+\theta_{b}+\theta_{c}=\pi$ we have
$\partial(a,a,a,b,c)=-\partial(a,a,b,c,a)-\partial(a,a,c,a,b).$
@@
Given that $\partial(a,b,c,d,e)=\partial(b,a,c,d,e)$, the only remaining unfriendly terms we have to treat are those of the form $\partial(b,a,a,b,c)$ and $\partial(b,a,a,c,b)$. We will treat the first one. The second one has the same treatment.
Thinking of our partial derivatives as operators on the function $\theta_{cde}$, we have the basic formula
$\frac{\partial}{\partial z_{c}}+\frac{\partial}{\partial z_{d}}+\frac{\partial}{\partial z_{e}}=0.$
Geometrically, the variation we are considering simply translates the triangle up or down along the $Z$-axis and changes none of the angles. Using this relation, we have
$\partial(b,a,a,b,c)=-\partial(b,b,a,b,c)-\partial(b,c,a,b,c).$
@@
In summary each partial derivative $\partial_{ij}\theta_{k}$ is a sum of $K$ friendly terms of the form $\partial(a,b,c,d,e)$ where $K\leq 24\times 2=48$. In view of this breakdown, it suffices to show $|\partial(a,b,c,d,e)|<10^{23}$ for all friendly terms.
5.4 The Final Estimates
We write vectors as $V=(u,v,w)$. We say that a pair $(V_{1},V_{2})$ of vectors is normalized if they have the form
$V_{1}=(0,v_{1},w_{1}),\hskip 30.0ptV_{2}=(u_{2},v_{2},w_{2}).$ (18)
We say that a torus $T^{\prime}$ is relevant if it corresponds to a point in the ball of radius $10^{-27}$ about the point describing our example $T$ in Equation 1. We say that a pair of normalized vectors is relevant if it is obtained from a triangle $(P_{0},P_{1},P_{2})$ in a relevant torus by the following procedure.
  • We set $V_{1}^{\prime}=P_{1}-P_{0}$ and $V_{2}^{\prime}=P_{2}-P_{0}$.
  • We rotate $(V_{1}^{\prime},V_{2}^{\prime})$ about the $Z$-axis to get a normalized pair $(V_{1},V_{2})$.
We make this construction for all permutations of these points. Thus each triangle in a relevant torus gives rise to $6$ pairs of relevant vectors.
Let $\vartheta(V_{1},V_{2})$ be the angle between $V_{1}$ and $V_{2}$. If we have an arbitrary pair of vectors we can rotate about the $Z$-axis so that they are normalized. This rotation does not change the geometry of the triangle defined by these vectors and it does not change the dependence of this geometry on the $Z$-coordinate. For this reason, each friendly term discussed in the previous section has the form
$\partial_{ij}\vartheta=\frac{\partial^{2}\vartheta}{\partial w_{i}\partial w_{j}},\hskip 30.0pt\vartheta=\arccos\bigg{(}\frac{V_{1}\cdot V_{2}}{\sqrt{(V_{1}\cdot V_{1})(V_{2}\cdot V_{2})}}\bigg{)}$ (19)
Let $g_{ij}=(\partial_{ij}\vartheta)^{2}$. This is a rational function which we compute explicitly below. Note also that $g_{12}=g_{21}$. So, to prove the Crude Bound Lemma, we just need to prove for all relevant pairs that
$g_{ij}<10^{46},\hskip 30.0pt(i,j)=(1,2),(1,1),(2,2).$

### v9

@@ 3.1 Robust Embedding
Remark: There is another approach we could have taken which involves computing the signs of the volumes of all the tetrahedra made from vertices of $T$. See [DS].
3.2 The Proof Modulo a Crude Bound
@@
Let $G:\mbox{{$R$}}^{3}\to\mbox{{$R$}}^{3}$ be a smooth map. Let $B_{r}(p)$ denote the ball of radius $r$ about $p$. We say that $G$ is $\lambda$-expansive on $B_{r}(p)$ if, for all unit vectors $V$ and all $q\in B_{r}(p)$, we have $\|dG_{q}(V)\|>2\lambda$ and $\angle(dF_{p}(V),dF_{q}(V))<\pi/3$.
@@
Proof: Let $O$ be the origin and let $B=B_{1}(O)$. By scaling and translation, it suffices to prove that $B\subset G(B)$ when $p=q=O$ and $\lambda=2$. Let $\alpha$ be any line segment connecting $O$ to a point on $\partial B$. Let $s\mapsto\alpha(s)$ be the unit speed parametrization of $\alpha$. Let $\beta=G(\alpha)$. The parametrization $s\mapsto\beta(s)$ has speed at least $2$. Since $\angle(\beta^{\prime}(t),\beta^{\prime}(0))<\pi/3$, the projection of $\beta$ onto the line through $\beta^{\prime}(0)$ moves with speed greater than $1$. But then $\|\beta(1)-\beta(0)\|>1$. This shows that $G(\partial B)$ lies entirely outside $B$. But now the angle condition implies that $G:\partial B\to\mbox{{$R$}}^{3}-B$ is homotopic to the identity map in $\mbox{{$R$}}^{3}-B$. Hence $G(\partial B)$ is the generator of the homology group $H_{2}(\mbox{{$R$}}^{3}-q)$ for any $q\in B$. This easily implies that $B\subset G(B)$. $\spadesuit$
Let $F(z_{0},z_{2},z_{3})=(\theta_{0},\theta_{2},\theta_{3})$ be the map from §2. Let $p\in\mbox{{$R$}}^{3}$ be the point that corresponds to our torus $T$ from Equation 1. We prove the following result in the next section.
@@
Throughout $B_{10^{-4}}(p)$ we have
$\bigg|\frac{\partial^{2}\theta_{k}}{\partial z_{i}\partial z_{j}}\bigg|<10^{9},\hskip 30.0pt\forall\ i,j,k\in\{1,2,3\}.$
Here is the main corollary.
@@
$F$ is $10^{-2}$-expansive on $B_{10^{-12}}(p)$.
Proof: Let $V$ be a unit vector. Define
$M=\left[\begin{matrix}-1.66&+1.46&-0.02\\ +1.46&+0.04&-1.13\\ -0.02&-1.13&+1.91\end{matrix}\right]$ (10)
$M$ is a symmetric and its eigenvalues all exceed $3/10$ in absolute value. Hence $\|M(V)\|>3/10$. We will settle for the weaker bound:
$\|M(V)\|>\frac{1}{10}.$ (11)
Define “error matrices” $E_{q}$ and $E^{\prime}_{q}$ so that
$dF_{q}=M+E_{q}=dF_{p}+E^{\prime}_{q},\hskip 30.0ptE_{q}=E_{p}+E_{q}^{\prime}.$ (12)
Comparing Equations 3 and 10 we see that
$\|E_{p}\|_{\infty}<\frac{1}{100}.$ (13)
We can join $p$ to $q$ by an axis-aligned path of length at most $\sqrt{3}\times 10^{-12}$. Hence, by the Crude Bound Lemma and integration:
$\|E^{\prime}_{q}\|_{\infty}<\sqrt{3}\times 10^{-12}\times 10^{9}<\frac{1}{450}\quad\Longrightarrow\quad\|E^{\prime}_{q}(V)\|<\frac{1}{150}.$ (14)
By Equation 13, Equation 14, and the triangle inequality,
$\|E_{q}\|_{\infty}\leq\|E_{p}\|_{\infty}+\|E^{\prime}_{q}\|_{\infty}<\frac{1}{100}+\frac{1}{450}<\frac{1}{75}\quad\Longrightarrow\quad\|E_{q}(V)\|<\frac{1}{25}.$ (15)
By Equation 11, Equation 15, and the triangle inequality,
$\|dF_{q}(V)\|\geq\|M(V)\|-\|E_{q}(V)\|>\frac{1}{10}-\frac{1}{25}>2\times\frac{1}{100}.$ (16)
By Equation 14, Equation 16, and some elementary geometry,
$\frac{\|E^{\prime}_{q}(V)\|}{\|dF_{p}(V)\|}<\frac{1/150}{1/50}=\frac{1}{3}<\frac{\sqrt{3}}{2}\quad\Longrightarrow\quad\angle(dF_{p}(V),dF_{q}(V))<\frac{\pi}{3}.$ (17)
Equations 16 and 17 give the desired bounds. $\spadesuit$
@@
Let $q=F(p)$ and let $q^{*}=(2\pi,2\pi,2\pi)$. We compute directly, in Mathematica, that $\|q-q_{*}\|<10^{-31}$. By the Surjectivity and Expansion Lemmas, $q^{*}=F(p^{*})$ with $\|p-p^{*}\|<10^{-29}$. The torus $T^{*}$ corresponding to $p^{*}$ is flat. Since $T$ is $10^{-4}$ robustly embedded, we see that $T^{*}$ is embedded. The dihedral angles of $T$ all lie in $005,\pi-.005]$, so the tiny change we make in going from $T$ to $T^{*}$ keeps $T^{*}$ a pup tent. This completes the proof of Theorem 1.2 modulo the Crude Bound Lemma.
Remark: All we used for our proof is that $T$ is $10^{-12}$ robustly embedded and that $\|q-q^{*}\|<10^{-14}$. One can get these weaker bounds from ordinary floating point operations in Java.
3.3 Proof of the Crude Bound Lemma
We keep the notation from above. Note that when we vary $(z_{1},z_{2},z_{3})$, we also vary $(z_{6},z_{5},z_{4})$ in a symmetric way. Thus,
$\frac{\partial\theta_{j}}{\partial z_{1}}=\partial_{1}\theta_{j}+\partial_{6}\theta_{j},\qquad\frac{\partial^{2}\theta_{j}}{\partial z_{1}\partial z_{2}}=(\partial_{12}+\partial_{62}+\partial_{15}+\partial_{65})\theta_{j}.$
$\partial_{i}\theta_{j}$ denotes the rate of change of $\theta_{j}$ when we vary only $z_{i}$ and $\partial_{ij}\theta_{k}$ denotes the rate of change of $\partial_{j}\theta_{k}$ when we vary only $z_{i}$. Here we are leaving the realm of $2$-fold symmetric configurations and just considering $8$ points in space.
@@ 3.3 Proof of the Crude Bound Lemma
Each $\theta_{j}$ is a $6$-term expression of terms of the form $\theta_{ijk}$, where $\theta_{ijk}$ is the angle of triangle $(i,j,k)$ at vertex $i$. Hence, $\partial_{ij}\theta_{j}$ is a $24$-term sum of expressions of the form $\partial_{ab}\theta_{cde}$. These expressions vanish when $\{a,b\}\not\subset\{c,d,e\}$. We call this expression friendly if $c\not\in\{a,b\}$, and otherwise unfriendly.
@@
Proof: Consider terms of $\partial_{aa}\theta_{abc}$. Using the relation $\theta_{a}+\theta_{b}+\theta_{c}=\pi$ we have
$\partial_{aa}\theta_{abc}+\partial_{aa}\theta_{bca}+\partial_{aa}\theta_{cab}=0.$
@@
Given that $\partial_{ab}=\partial_{ba}$, the only remaining unfriendly terms we have to treat are those of the form $\partial_{ba}\theta_{abc}$ and $\partial_{ba}\theta_{acb}$. We will treat the first one. The second one has the same treatment.
We also have the relation $(\partial_{a}+\partial_{b}+\partial_{c})\theta_{abc}=0$. Geometrically, the variation we are considering simply translates the triangle up or down along the $Z$-axis and changes none of the angles. Applying $\partial_{b}$ to this relation, we have
$\partial_{ba}\theta_{abc}=-\partial_{bb}\theta_{abc}-\partial_{bc}\theta_{abc}.$
@@
In summary, each partial derivative in the Crude Bound Lemma is a sum at most $48$ friendly term. To prove the Crude Bound Lemma it suffices to show that each friendly term is less than $10^{7}$ in absolute value throughout the ball $B_{10^{-4}}(p)$.
We say that a torus is relevant if it corresponds to a point in our tiny ball $B$. We say that a pair $(V_{1},V_{2})$ of vectors is relevant if it is obtained from a triangle $(P_{0},P_{1},P_{2})$ of a relevant torus by setting $V_{1}=P_{1}-P_{0}$ and $V_{2}=P_{2}-P_{0}$. Each triangle in a relevant torus gives rise to $6$ pairs of relevant vectors.
Let $\vartheta(V_{1},V_{2})$ be the angle between $V_{1}$ and $V_{2}$. Each friendly term has the form
$\frac{\partial^{2}\vartheta}{\partial w_{i}\partial w_{j}},\hskip 30.0pt\vartheta=\arccos\bigg(\frac{V_{1}\cdot V_{2}}{\sqrt{(V_{1}\cdot V_{1})(V_{2}\cdot V_{2})}}\bigg)$ (18)
The square $g_{ij}$ of the expression on the left side of Equation 18 is a rational function. To prove the Crude Bound Lemma, we just need to prove for all relevant pairs and all $(i,j)$ that that $g_{ij}<10^{14}$.